Status Report

SOHO Weekly Meeting Minutes – Week 08, February 18, 2002

By SpaceRef Editor
February 18, 2002
Filed under , ,

1. Announcements
* Since Monday is a holiday, there will be no daily meeting.
A student SOC will be in the EOF for part of the day.

* CTOF recovery will begin at 16 UT Tuesday and will likely continue
on Wednesday. Expect at least 10 minutes NRT pause Tuesday 16 UT.

* CMS Testing for an hour, probably on Thursday morning, during downlink
only station.

* SPWG on Friday at 2pm in the EAF

* Two members of the House Appropriations Committee staff are
currently scheduled to visit the SOHO EOF at 1:00 PM on
Friday, March 1, for 20 – 25 minutes.

2. Operations Constraints
* Commanding schedule for the week:
Monday – D66 until 16:30, D27 after 19 UT.
Tuesday – D66 until 18:30 then D27.
Wedesday – D66 until 18:30 then D27.
Thursday – D/L only in the morning then D27 after 17:30 UT.
Friday – D/L only until 19 UT then D16.
Saturday – Command on D27 all day.
Sunday – Command on D27 all day.

* UVCS Star Observations on Thursday

3. Coordinated Observations
M Feb 18 (W08) President’s Day
T Feb 19
W Feb 20
T Feb 21 UVCS Star Observation of 38 Aqr

F Feb 22 SPWG at 2pm
S Feb 23
S Feb 24

Other activities for Week 08:
* Feb 9-26 UVCS Observations of a High Latitude Southern Coronal Hole
* Feb 20-Mar 7 UVCS Streamer and Slow Solar Wind Study

Targets of Opportunity (TOO):
* Feb 19-28 JOP152 Multi-wavelength Observations of Large-scale
Shock Waves on the Sun, CDS/EIT/TRACE,
POC: Josef Khan, Terry Kucera

Submodes: Submode 6

Planners for Week 08:
CDS — Alex Young MDI — Sarah Gregory
EIT — Joe Gurman SUMER — not observing
LASCO — Ops Team UVCS — Steve Cranmer/Leonard Strachan

Relevant Notes
CDS: JOP152, usual weekly studies, possible HESSI support.

UVCS: Coronal hole observations and synoptics. Flat field
calibration on Wednesday, star observations on Thursday.

LASCO: C2/C3 synoptics.

EIT: 195 CME watch and synoptics, supporting JOP152, supporting

MDI: Full disk magnetograms and doppergrams for continuous

TRACE: AR tracking study, switching to 195A during the week.

SpaceRef staff editor.